# Thought Anchors: Which LLM Reasoning Steps Matter?

### v3

@@ 4.1 The identification of receiver heads
We refer to attention heads that narrow attention toward specific sentences as “receiver heads”. These heads are more common in later layers (Appendix˜H). To formally assess the existence of receiver heads, we tested whether some attention heads consistently operate in this role. Specifically, we measured the split-half reliability of heads’ kurtosis scores. We found a strong head-by-head correlation ($r$ = .67) between kurtosis scores computed for half of the problems with kurtosis scores for the other half of problems. Thus, some attention heads consistently operate as receiver heads, albeit with some heterogeneity across responses in which heads narrow attention most.
@@
Figure 4: A. Lines show the vertical attention scores for each sentence by the 40 different heads in layer 36. Head 6 has been highlighted as a receiver head, and its corresponding attention weight matrix is shown for reference. Its prominent spikes cause the distribution to have a high kurtosis. B. Histogram of these kurtosis values across all attention heads, averaged across all reasoning traces.
@@
Receiver heads usually direct attention toward the same sentences. Among the 16 heads with the highest kurtoses, we computed the sentence-by-sentence correlation between the vertical-attention scores for each pair of heads; calculated separately for each reasoning trace, then averaged. This produced an large correlation (mean $r$ = .60). Thus, receiver heads generally attend the same sentences, albeit with some heterogeneity (this correlation among randomly selected heads is $r$ = .45). This convergence is consistent with the existence of sentence importance, which these heads identify.
@@
Attentional narrowing toward particular sentences may be a feature specifically of reasoning models that enhances their performance. Comparing R1-Distill-Qwen-14B (reasoning) and Qwen-14B (base) suggests that the reasoning model’s receiver heads will narrow attention toward singular sentences to a greater degree (appendix˜I). Furthermore, ablating receiver heads leads to a greater reduction in accuracy than ablating self-attention heads at random (appendix˜J). Altogether, these findings are consistent with receiver heads and thought-anchor sentences playing special roles in reasoning.
4.2 Links to resampling importance and sentence types
Plan generation, uncertainty management, and self checking sentences consistently receive the most attention via receiver heads (Figure 5), whereas active computation sentences receive minimal attention. Further consistent with this, sentences receiving high receiver-head attention tend to also exert a larger effect on downstream sentences per the resampling method (mean Spearman $\rho=.22$; 95% CI: [.17, .30]; detailed further in Appendix˜K). These findings are hence consistent with reasoning traces being structured around high-level sentences that initiate computations that may link high-level statements but may minimally impact the overall reasoning trajectory.
@@ 4.2 Links to resampling importance and sentence types
Figure 5: The boxplot shows the average top-16 receiver-head score for each sentence type. The boxes correspond to the interquartile range across different reasoning traces.
5 Attention suppression
@@ 5.1 Approach
A natural concern with examining attention weights is that they do not measure causality. Further, receiver heads are not necessarily suited for identifying sentences’ links to individual future sentences. Hence, we next focused on causal attribution for sentence-sentence dependencies. The sentence-sentence analyses based on resampling also examined such dependencies but had limited precision for mapping logical connections, as counterfactual importance represents the total effect (direct and indirect) of one sentence on another. The present method instead aims to isolate the direct effects between sentences, which we expect to permit more precise modeling of logical connections.
We suppress all attention (all layers and heads) toward a given sentence and examine the impact on future sentences. The impact is defined as the KL divergence between a token’s logits and its baseline logits without suppression. The effect on a future sentence is calculated as the average of its constituent token log-KL divergences. This technique assumes (i) token logits accurately capture a sentence’s semantic content and (ii) suppressing attention does not problematically induce out-of-distribution behavior. Hence, we must validate correspondence between the attention-suppression measure and resampling measure, which, although less precise, does not hinge on these assumptions.
5.2 Correlations with the resampling-based importance matrix
The attention-suppression matrix values correlate with those of the resampling-method matrix. Specifically, the two matrices were positively correlated for 19 of the 20 examined reasoning traces (mean: Spearman $\rho=.20$, 95% CI: [.12, .27]). This association is stronger when considering only cases fewer than five sentences apart in the reasoning trace, which may better track direct rather than indirect effects represented by the resampling method (mean: $\rho=.34$ [.27, .40]). The magnitudes of these correlations are substantial, given that measures measure partially different aspects of causality and the resampling measure itself contains stochastic noise. Hence, these results give weight to the validity of the resampling approach, whose precision we leverage for the forthcoming case study.
6 Case study
The three presented techniques cover different aspects of attribution within a reasoning trace. We will now illustrate the utility and the complementary nature of these techniques by applying them to the model’s response for one problem: “When the base-16 number $66666_{16}$ is written in base 2, how many base-2 digits (bits) does it have?” (Figure 6) (see Appendix˜L for the full transcript).
6.1 Resampling
The model approaches this problem by first considering that $66666_{16}$ contains five base-16 digits and any base-16 digit can be represented with four base-2 digits. Following this logic, the CoT starts by pursuing the answer “20 bits”. However, this initial answer overlooks that $6_{16}$ is $110_{2}$ (i.e., not $0110_{2}$ due to the leading zero), so the correct answer is actually “19 bits”. In sentence 13, the model pivots toward this correct answer by initiating computations to convert $66666_{16}$ to decimal and then binary.
The resampling method (section˜3) reveals this initial poor trajectory and pivot upwards (shown previously in Figure 2A). Specifically, from sentences 6-12, expected accuracy steadily declines, but sentence 13 causes counterfactual accuracy to drastically increase. The pivotal role of 13 is notably missed if one instead evaluates accuracy by forcing the model to give a response immediately, as some prior studies have done, which instead yields 0
6.2 Receiver heads
6.3 Attention suppression
@@ 6.4 Open source interface
We released an open source interface (thought-anchors.com) for analyzing and visualizing reasoning traces. The reasoning trace can be visualized as an annotated directed acyclic graph using the importance metrics from Section˜3, which may benefit interpretability and reasoning model debugging. We show our proof-of-concept interface in Figure 1C, where important sentences are represented by larger nodes and sentences that causally affect each other are connected with dashed gray lines. Hovering over a node lets users inspect sentence properties or compare alternative rollouts.
@@ 7 Related work
Reasoning advances in LLMs.
Chain-of-thought reasoning (Wei et al., 2023; Nye et al., 2021; Reynolds and McDonell, 2021) has driven significant capabilities improvements in large language models. State-of-the-art performance on complex tasks comes from “thinking models” like DeepSeek-R1 (DeepSeek, 2025), OpenAI’s o1 (OpenAI, 2024), and Claude 3.7 (Anthropic, 2025) are explicitly optimized for this with RLVR(Yue et al., 2025). These increased reasoning capabilities introduce novel safety challenges, including potential unfaithfulness in the reasoning itself (Lanham et al., 2023b). While current alignment evaluation methods often rely on examining reasoning traces (Baker et al., 2025; Shah et al., 2025), we cannot rely on the model’s reasoning being faithful to the internal reasoning(Chen et al., 2025; Turpin et al., 2023; Arcuschin et al., 2025), creating a need for more mechanistic analyses.
Importance of individual steps. Many CoT sentences can be deleted or rewritten with limited impact on task accuracy (Wang et al., 2023; Madaan and Yazdanbakhsh, 2022). Prior attribution methods suggest that a subset of steps disproportionately drive the final answer: Shapley values (Gao, 2023), ROSCOE metrics (Golovneva et al., 2023), gradient-based scores (Wu et al., 2023), and resampling at fork tokens (Bigelow et al., 2024). Complementing these, our approach provides a more principled framework for understanding how reasoning traces are constructed around key sentences.
White-box analyses of reasoning. Mechanistic studies locate specialised components that propagate information across steps: iteration or receiver heads (Cabannes et al., 2024), small sets of heads necessary for CoT performance (Zheng et al., 2024), steerable activation directions for planning and backtracking (Venhoff et al., 2025), and deduction circuits in symbolic tasks (Brinkmann et al., 2024). Other work identifies phase shifts separating pre-training priors from contextual reasoning (Dutta et al., 2024). These findings motivate our focus on sentence-level structure rather than token-level saliency. However, they do not provide general methods for understanding the causal interdependence between sentences.
@@ 8 Discussion and Limitations
We view this as preliminary work. Our analyses require refinement to fully grapple with how downstream sentences may be overdetermined by different trajectories in a reasoning trace or independent sufficient causes. Additionally, we do not formally examine the role of error correction. Our receiver-head analyses are confounded by a sentence’s position in the reasoning trace (see Appendix˜M). Further, our attention-suppression work is limited because it effectively requires the model to process out-of-distribution information.
Nonetheless, we believe that we have demonstrated that our metrics are an advance on prior work, interrupting models and forcing final answers. The surprising degree of shared structure we have found across our three methods illustrates the potential value of future work in this area and points to the possibility of more powerful interpretability techniques to come.
@@ Acknowledgments
This work was conducted as part of the ML Alignment & Theory Scholars (MATS) Program. We would like to thank Iván Arcuschin, Constantin Venhoff, and Samuel Marks for helpful discussions and feedback. We particularly thank Stefan Heimersheim for his valuable feedback and insightful suggestions, including ideas for experimental approaches that helped strengthen our analysis and contributed to the clarity of our presentation. We also thank members of Neel Nanda’s MATS stream for engaging brainstorming sessions, thoughtful questions during our presentations, and ongoing discussions that helped shape our approach.
@@
Appendix A Sentence taxonomy
Building on top of the framework presented by [Venhoff et al., 2025], we developed a taxonomy consisting of eight distinct sentence categories that capture reasoning functions in mathematical problem-solving. Each category represents a specific cognitive operation. The functions and examples for each category are given in Table 1. Notably, the uncertainty management category includes backtracking sentences.
Category Function Examples
Appendix B Prompt information
Function Tags:
Dependencies:
- If there’s no clear dependency use an empty list: []
Important Notes:
- Make sure to include all dependencies for each sentence.
- Include both long-range and short-range dependencies.
- Do NOT forget about long-range dependencies.
- Try to be as comprehensive as possible.
- Make sure there is a path from earlier sentences to the final answer.
Output Format:
Return a dictionary with one entry per sentence, where each entry has:
- the sentence index (as the key, converted to a string),
- a dictionary with:
- "function_tags": list of tag strings
- "depends_on": list of sentence indices, converted to strings
Here is the expected format:
"1": {
"function_tags": ["problem_setup"],
"depends_on": [""]
"4": {
"function_tags": ["plan_generation"],
"depends_on": ["3"]
"5": {
"function_tags": ["fact_retrieval"],
"depends_on": []
"9": {
"function_tags": ["active_computation"],
"depends_on": ["4", "5"]
"24": {
"function_tags": ["uncertainty_management"],
"depends_on": ["23"]
"32": {
"function_tags": ["final_answer_emission"],
"depends_on": ["9, "30", "32"]
Here is the math problem:
<PROBLEM>
Here is the full chain-of-thought, broken into sentences:
<SENTENCES>
Now label each sentence with function tags and dependencies.
Appendix C Sentence category probing
Appendix D Other reasoning model
D.1 Measuring counterfactual influence
To assess the generalizability of our counterfactual importance findings, we replicated our resampling methodology on R1-Distill-Llama-8B, applying the same experimental parameters (e.g., temperature = $0.6$ and top-p = $0.95$) used for R1-Distill-Qwen-14B. We collected 100 rollouts for 10 correct and 10 incorrect base solutions using the identical question set described in Section 2.

### v4

@@ 4.1 The identification of receiver heads
We refer to attention heads that narrow attention toward specific sentences as “receiver heads”. These heads are more common in later layers (Appendix˜J). To formally assess the existence of receiver heads, we tested whether some attention heads consistently operate in this role by measuring the split-half reliability of heads’ kurtosis scores. We found a strong head-by-head correlation ($r$ = .84) between kurtosis scores computed for half of the problems with kurtosis scores for the other half of problems. Thus, some attention heads consistently operate as receiver heads, albeit with some heterogeneity across responses in which heads narrow attention most.
@@
Figure 4: A. Lines show the vertical attention scores for each sentence by the 40 different heads in layer 36. Head 6 has been highlighted as a receiver head, and its corresponding attention weight matrix is shown for reference. Its prominent spikes cause the distribution to have a high kurtosis. B. Histogram of these kurtosis values across all attention heads, median across all reasoning traces.
@@
Receiver heads usually direct attention toward the same sentences. Among the 16 heads with the highest kurtoses, we computed the sentence-by-sentence correlation between the vertical-attention scores for each pair of heads; calculated separately for each reasoning trace, then averaged. This produced an large correlation (mean $r$ = .56). Thus, receiver heads generally attend the same sentences (for reference, the average correlation among any heads is $r$ = .35). This convergence across receiver heads is consistent with the existence of sentence importance, which these heads identify.
@@
Attentional narrowing toward particular sentences may be a feature specifically of reasoning models that enhances their performance. Comparing R1-Distill-Qwen-14B (reasoning) and Qwen-14B (base) suggests that the reasoning model’s receiver heads will narrow attention toward singular sentences to a greater degree (Appendix˜K). Furthermore, ablating receiver heads leads to a greater reduction in accuracy than ablating self-attention heads at random (Appendix˜L). Altogether, these findings are consistent with receiver heads and thought-anchor sentences playing special roles in reasoning.
4.2 Links to counterfactual importance and sentence types
Plan generation and uncertainty management sentences consistently receive the most attention via receiver heads (Figure 5), whereas active computation sentences receive relatively minimal attention ($t$s > 4.0, $p$s < .001 per paired t-tests comparing the mean receiver-head score for the former two versus the later two categories). These findings demonstrate a parallel between the receiver head findings here and the earlier results on the sentence types yielding the highest counterfactual importance.
@@ 4.2 Links to counterfactual importance and sentence types
Figure 5: The boxplot shows the average top-32 receiver-head score for each sentence type. The boxes correspond to the interquartile range across different reasoning traces.
5 Sentence-sentence causal links
@@ 5.1 Approach
We next focused on how key sentences influence specific subsequent sentences. We used a sentence-masking strategy, suppressing all attention (all heads) to a given sentence and examining how this impacts future sentences, measured using the KL divergence between token logits with or without masking. The overall effect on a future sentence is the average of its token log-KL divergences. We normalize this score by subtracting the latter sentence’s average causal effect from all prior sentences. Suppressing attention is mostly equivalent to omitting a sentence from a CoT, only differing in positional embeddings; computing the causal graph as so can be done with LLM APIs that return logits without exposing attention. Appendix˜M provides pseudocode for generating the causal graph.
Our masking approach assumes (i) token logits capture a sentence’s semantic content and (ii) masking sentences does not problematically induce out-of-distribution behavior. We evaluated these assumptions by correlating the sentence-sentence scores with those from an alternative strategy based on our counterfactual resampling method, which assesses how resampling $S_{i}$ with $T_{i}$ ($S\not\approx T$) influences the likelihood of $S_{j}$ appearing. This measure positively correlates with the scores from the masking- & logits-based strategy (detailed further in appendix˜N), suggesting that logits indeed track semantics despite simulating out-of-distribution behavior. We continue with the sentence-masking approach because it requires $\sim$100x less compute than resampling, increasing scalability.
5.2 Case study
We begin with a small-scale investigation to provide intuition for our sentence-sentence measure and motivate more systematic tests. We continue our initial case study (Section˜2.4), but here, we focus on three local maxima in the sentence-masking graph (Figure 6), which align closely with the sentences implicated as important by receiver-heads (see further details on the case study in Section˜C.3):
  • (Sentences: 12 → 43) After suggesting the answer “20 bits”, the model decides to begin verifying it (Sentence 12). Verification leads to a different solution, “19 bits” (Sentence 43). Between these key sentences, most of the intermediate text is performing arithmetic.
  • (Sentences: 44 → 65) Noticing the discrepancy (Sentence 44), the model decides to check its calculations. It finds that they are correct, and the discrepancy remains (Sentence 65).
  • (Sentences: 12 → 66) The model realizes that its initial suspicion about leading zeroes (Sentence 12) is justified and states that this is the reason for the discrepancy (Sentence 66).
These connections point to an interpretable scaffold reflecting computations on the pursuit of intermediate results, the execution of self-correction subroutines, and the synthesis of prior statements.
Figure 6: For the correct-answer CoT of Problem #4682, the matrix shows the effect of masking one sentence (x-axis) on a future sentence’s logits (y-axis). Darker colors indicate higher values.
@@ 5.3 Open source interface
We released an open source interface (thought-anchors.com) for visualizing reasoning traces and comparing alternative rollouts. We show our proof-of-concept interface in Figure 1C, where important sentences are represented by larger nodes and causal connections between sentences are shown with dashed gray lines. The tool aims to benefit interpretability and unwanted behavior debugging.
6 Systematic differences in sentence-sentence relationships
The case study demonstrates that pivotal moments within a CoT – e.g., where a conclusion is posed, where a decision is made about the next steps to answering a question, or where a discrepancy with earlier information is identified – are sensitive to historic information in the CoT and/or exert impacts far downstream. We next investigated how causal graphs may shed light on questions about reasoning in LLMs generally. Specifically, we ask: How many examining sentence-sentence links shed light on model confidence during reasoning? Relatedly, why do some problem domains like mathematics display stronger uplift in reasoning compared to non-reasoning models?
We hypothesize that strong linkages between nearby sentences reflect a coherent logical flow and well-formed plan, so each sentence causes the next, whereas distant linkages reflect uncertainty and backtracking. Despite occasional long-range connections, we further hypothesize that successful mathematical CoTs are specifically characterized by strong close links between sequential sentences, whereby planning statements sharply structure the CoT. Domains related to mathematics may uniquely lend themselves to such firmly-structured reasoning, whereas CoTs for other topics (e.g., history or biology) may solve problems principally by scanning a wide latent space in a less structured fashion.
6.1 Methods
We pivoted to analyzing MMLU problems (Hendrycks et al., 2020), so that we could contrast problem domains. We also switched to Qwen3-30b-a3b, so that we could leverage a serverless LLM provider that outputs token logits, which allowed scaling up our analysis to thousands of CoTs. We ran Qwen3-30b-a3b in non-reasoning mode on all 15,638 MMLU questions to identify challenging problems where non-reasoning accuracy is under 50% (per answer logits). This corresponds to 3,651 problems, and for 2,492 of these questions, the model answers correctly when using reasoning at least once across ten passes. We computed each correct CoT’s causal graph ($M_{sentences}$ = 90.1).
We compared graphs on the strength of their causal links at different distances between sentences. We specifically computed the mean attention-suppression effect at distance $k$ for each graph ($m\times m$ sentences) for all $k\leq\frac{m}{2}$. This corresponds to the mean of a matrix’s $k$-th subdiagonal. We consider subdiagonals only up to $\frac{m}{2}$ to reduce noise by ensuring that the mean is computed among an adequate number of elements (e.g., the $m$-th subdiagonal would be just the single bottom-leftmost element).
6.2 Results
The distance of causal effects tracks question difficulty. Computing correlations within-subject, we find that questions with high average accuracy elicit CoTs with stronger close-range links and weaker long-range links (Figure 7A). In addition, subjects where average accuracy is high overall tend to produce CoTs with stronger close links ($r=.44,p<.001$; Figure 7B) and weaker long links (r=.54,p<.001r=-.54,p<.001 Figure 7C). The strongest levels of accuracy were seen in problems requiring mathematical thinking (e.g., mathematics & physics). As hypothesized, these areas also yielded CoTs with stronger close-range connections and weaker long-range connections (two-sample t-test $|t|s>10,ps<.001$; Figure 7D). Although these analyses do not model plan generation and uncertainty management (thought anchors) sentences directly, the present findings speak to their potential structural roles and overall shed light on the nature of successful reasoning.
Figure 7: A. For each distance $k$, we computed the correlation between a question’s average $k$-distance causal effect in one CoT and the question’s mean reasoning accuracy across ten CoTs. B. & C. Scatterplot shows each subject’s average close-range ($k$ = 1-2) and long-range ($k$ = 16-64) was plotted against its average reasoning accuracy. D. Box-plots showing the spread of average close-range and long-range causal effects for different question domains; each point represents one CoT, and black circles represent means.
@@ 7 Related work
Reasoning advances and unfaithfulness in LLMs.
CoT reasoning, optimized using reinforcement learning, has driven major capabilities improvements in large language models (Wei et al., 2023; Nye et al., 2021; Reynolds & McDonell, 2021). This reasoning paradigm introduces novel safety challenges. Experiments inducing unfaithful reasoning have led some to raise concerns about the interpretability of CoT text (Lanham et al., 2023b; Chen et al., 2025), although others have argued that CoT text generally is a meaningful representation, particularly for difficult tasks (Korbak et al., 2025). By showing how sentence types, categorized based on their text, differ in their resampling and receiver-head importance, our findings endorse the meaningfulness and interpretability of CoT text.
Importance of individual steps. A variety of prior techniques that can be used for CoT interpretability have been developed, and these likewise have suggested that a subset of steps disproportionately drive the final answer – e.g., findings based on Shapley values (Gao, 2023), ROSCOE metrics (Golovneva et al., 2023), gradient-based scores (Wu et al., 2023), and resampling at fork tokens (Bigelow et al., 2024). Complementing these, our approach provides a more principled framework for understanding how reasoning traces are constructed around key sentences.
@@ 8 Discussion and Limitations
While some research raises valid concerns that CoT text can be unfaithful to the model’s underlying computation (Lanham et al., 2023b; Chen et al., 2025), our results show CoT text is mechanistically relevant and interpretable. For example, sentences categorized as plan generation and uncertainty management consistently exhibit higher counterfactual importance in our resampling analyses and receive more focused attention from receiver heads. This demonstrates a link between what a sentence says and its functional role in the computation, and this type of correspondence supports arguments on the value of CoT legibility (Korbak et al., 2025).
A primary limitation of our resampling approach is its computational cost. In this work, we resampled 100 times per sentence to achieve fairly precise estimates (in terms of final-answer accuracy, 95% CI corresponds to at worst $\pm 10\%$), which was sufficient to identify pivotal moments in the case study (Figure 2A). However, for analyses focusing on aggregate patterns across many CoTs, fewer resamples for any one CoT may suffice. Additionally, future work could develop adaptive resampling strategies that allocate more computational budget to potentially pivotal moments in the trace, maximizing precision while minimizing cost.
We view this as preliminary work. Our analyses require refinement to fully grapple with how downstream sentences may be overdetermined by different trajectories in a reasoning trace or independent sufficient causes. Additionally, we do not formally examine the role of error correction. Our receiver-head analyses are confounded by a sentence’s position in the reasoning trace (see Section˜C.4). Further, our attention-suppression work is limited because it effectively requires the model to process out-of-distribution information.
Despite these limitations, we believe that we have demonstrated that our metrics are an advance on prior work, interrupting models and forcing final answers. The surprising degree of shared structure we have found across our three methods illustrates the potential value of future work in this area and points to the possibility of more powerful interpretability techniques to come.
9 Reproducibility Statement
To ensure the reproducibility of our work, we provide comprehensive implementation details, code, and experimental specifications. Our code is publicly available at https://github.com/interp-reasoning/thought-anchors/, which includes all scripts for black-box resampling, receiver head analysis, and attention suppression experiments. We also provide an interactive visualization tool at https://www.thought-anchors.com for exploring reasoning traces and sentence-level causal dependencies. The complete prompt used for sentence taxonomy labeling is provided in Appendix˜E, including detailed instructions for function tags and dependency annotations. Our experimental setup uses DeepSeek R1-Distill-Qwen-14B (48 layers) with temperature 0.6 and top-p 0.95, tested on the MATH dataset (Hendrycks et al., 2020) focusing on problems with 25-75% solution rates. We specify exact hyperparameters including 100 rollouts per sentence for counterfactual resampling, cosine similarity threshold of 0.8 (median value) using all-MiniLM-L6-v2 embeddings (Appendix˜G), and identification of receiver heads via kurtosis scores of attention distributions. The sentence-sentence causal masking methodology is fully detailed in Section˜5, with validation through correlation with resampling-based measures (Appendix˜N). For MMLU experiments in Section˜6, we used Qwen3-30b-a3b on 2,492 problems where non-reasoning accuracy is below 50%, computing causal graphs for correct CoTs. Additional reproducibility details include: full case study transcript (Appendix˜C), sentence category distributions (Appendix˜D), receiver head ablation procedures with 128/256/512 heads (Appendix˜L), and cross-model validation on R1-Distill-Llama-8B (Appendix˜B). All models used are publicly available, and we provide pseudocode for the sentence-to-sentence importance calculation in Appendix˜N.
@@ Acknowledgments
This work was conducted as part of the ML Alignment & Theory Scholars (MATS) Program. We would like to thank Iván Arcuschin, Constantin Venhoff, and Samuel Marks for helpful discussions and feedback. We particularly thank Stefan Heimersheim for his valuable feedback and insightful suggestions, including ideas for experimental approaches that helped strengthen our analysis and contributed to the clarity of our presentation. We also thank members of Neel Nanda’s MATS stream for engaging in brain-storming sessions, thoughtful questions during our presentations, and ongoing discussions that helped shape our approach.
@@
Appendix A Evaluating importance (KL) while smoothing
The identified link between sentence’s category and its forced-answer or counterfactual importances were also measured while smoothing the final-answer distribution associated with each sentence. Smoothing was performed when computing the KL divergence between the two distribution and constitutes replacing the $\epsilon=10^{-9}$ term (originally used to avoid division by zero) with $\alpha=1.0$ (Laplace smoothing) or $\alpha=0.5$ (smoothing with Jeffrey’s prior).
Let $p(A^{\prime}_{S_{i}})$ and $p(A_{S_{i}})$ be the empirical distributions over a set of $K$ possible final answers, $\mathcal{A}$, derived from $N$ rollouts (e.g., $N=100$). Let $C^{\prime}_{S_{i}}(a)$ and $C_{S_{i}}(a)$ be the observed counts for a specific answer $a\in\mathcal{A}$ in the intervention and base conditions, respectively, such that $\sum_{a\in\mathcal{A}}C^{\prime}_{S_{i}}(a)=N$. Additive smoothing with a parameter $\alpha$ is applied to derive smoothed probabilities, $p_{\alpha}$ and $q_{\alpha}$, from these counts:
$p_{\alpha}(a)=\frac{C^{\prime}_{S_{i}}(a)+\alpha}{N+K\alpha}\quad\text{and}\quad q_{\alpha}(a)=\frac{C_{S_{i}}(a)+\alpha}{N+K\alpha}$
The smoothed KL divergence, $D_{\text{KL}}^{\alpha}$, is then computed using these non-zero probabilities:
$D_{\text{KL}}^{\alpha}[p(A^{\prime}_{S_{i}})||p(A_{S_{i}})]=\sum_{a\in\mathcal{A}}p_{\alpha}(a)\log\left(\frac{p_{\alpha}(a)}{q_{\alpha}(a)}\right)$
This method replaces the use of a small $\epsilon$ floor. The smoothing parameters used are $\alpha=1.0$ (Laplace smoothing) and $\alpha=0.5$ (Jeffreys prior).
With either level of smoothing, the same patterns linking importance and sentence category emerge as initially reported without smoothing (Figure 3). Specifically, active computation sentences yield higher forced answer importance than plan generation and uncertainty management, but the reverse is true when examining counterfactual importance based on the resampling method (Figure 8).
Figure 8: This is a variant of Figure 3, now performed with smoothing. Smoothing was performed using either (A) $\alpha=1.0$, Laplace smoothing, or (B) $\alpha=0.5$, Jeffreys prior.
Appendix B Generalizing to an alternative reasoning model
B.1 Measuring counterfactual influence
To assess the generalizability of our counterfactual importance findings, we replicated our resampling methodology on R1-Distill-Llama-8B, applying the same experimental parameters (e.g., temperature = $0.6$ and top-p = $0.95$) used for R1-Distill-Qwen-14B. We collected 100 rollouts for 20 correct and 20 incorrect base solutions using the identical question set described in Section 2.